OpenStudy (anonymous):

what is the value of y in the equation 4y-2(1-y)=-44

OpenStudy (tkhunny):

Have you considered the distributive property? -2(1-y) = -2+2y

OpenStudy (anonymous):

ok so first you have to multiply out your brackets which is -2(1-y)= -2+2y so you get 4y-2+2y=-44 so all you have to do is get y alone 4y+2y=-44+2 6y=-42 y=-42/6 y=-7